A car comes to a bridge during a storm and finds the bridge washed out. The driver must get to the other side, so he decides to try leaping it with his car. The side the car is on is 21.9m

above the river, whereas the opposite side is a mere 1.9m

above the river. The river itself is a raging torrent 53.0m

wide.

Part A

How fast should the car be traveling just as it leaves the cliff in order to just clear the river and land safely on the opposite side?

Part B

What is the speed of the car just before it lands safely on the other side?
Answer #1

A)

In vertical direction:
d = (21.9 - 1.9) = 20.0 m
a = 9.8 m/s^2
vi=0 m/s
use:
d = vi*t + 0.5*a*t^2
20 = 0 + 0.5*9.8*t^2
t= 2.02 s

In horizontal direction,
d = 53.0 m
t = 2.02 s
vx = s/t = 53.0/2.02 = 26.2 m/s
Answer: 26.2 m/s
B)
Final velocity in horizontal direction will be same.
vfx = 26.2 m/s

In vertical direction:
vf= vi+a*t
= 0 +9.8*2.02
=19.8 m/s
so,
vfy = 19.8 m/s

speed= sqrt (vfx^2 + vfy^2)
= sqrt (26.2^2 + 19.8^2 )
= 32.8 m/s
Answer: 32.8 m/s